mikewwe13:

The number of students enrolled in a new course as a function of time can be represented by the function. f(x) = (4)^x − 1 What is the average increase in the number of students enrolled per hour between hours 2 and 4? Enter your answer in the box. _______________

mikewwe13:

@Vocaloid

celticcat:

number enrolled after 2 hours = 4^2 - 1 = 15 after 4 hours it is 4^4 - 1 = 255 so the increase in 2 hours is 255 - 15 = 240 required average is 240 / 2 = 120